3 - 12 Let A = Construct a 2x2 matrix B such that AB is the zero matrix. Use two -4 16 different nonzero columns for B. B=
Answer #1

3 12 A -4 16 Let B: [ [] ] * [3][4] [3] 0 0 AB 0 0 3a -126 3C -124 -4a+16b -4C +6d 3q -12b = 0 3c-12d =0 - 44 +16 b = 0 -4c +